# Rigidity transition in polydisperse shear-thickening suspensions

### v1

@@ IV.2 Finite size scaling
Figure 7: Finite size scaling of the correlation length $\xi$ and susceptibility $\chi_{\text{rig}}$ for polydisperse suspensions with number of particles $N=1000$, $2000$ and $3000$, $\alpha=0.133$ at stress $\sigma=100$. Collapse of data for different system sizes are observed with scaling exponents $\nu=1.64$ and $\gamma=7/4$.
@@ IV.3 Effects of polydispersity index and particle stiffness on $\phi_{c}$
According to the present model, the frictional contacts between the particles depend on the spring constants for the normal and tangential components of the contact force. While in this study we have considered the particles to be non-deformable rigid spheres, the stiffness of the particles may still be correlated to the stiffness of the springs at the frictional contact. This section discusses the effects of polydispersity index and spring constants on the critical packing fraction and shear jamming. As presented in Fig. 8 (a), for a given value of stress ($\sigma$) and spring constant ($k_{n}$) the critical packing fraction $\phi_{c}$ varies non-monotonically with polydispersity index $\alpha$ and goes through a minima. The same trend is observed for the shear jamming fraction $\phi_{J}^{\mu}$, as shown in Fig. 8 (b). However, it is to be noted that for the same value of $k_{n}$(=2000=2000), while $\phi_{c}$ is smaller for the larger stress ($\sigma=100$) at all polydispersity indices (as shown in Fig. 8 (a)), $\phi_{J}^{\mu}$ shows a opposite behaviour where it is higher for $\sigma=100$ as compared to $\sigma=25$. Interestingly, by proportionally lowering the value of $k_{n}$ to 500 for $\sigma=25$, both $\phi_{J}^{\mu}$ and $\phi_{c}$ are found to be higher for $\sigma=25$ as compared to $\sigma=100$ at all values of $\alpha$ (except for very small $\alpha$). Thus with decreasing spring stiffness, both $\phi_{J}^{\mu}$ and $\phi_{c}$ increases at any given stress suggesting the suspensions to become more flowable with softer particle contacts. This can be correlated to increasing flowability of softer deformable particles which could easily squeeze past each other under shear flow as compared to more rigid particles. Moreover, the similar variation of $\phi_{J}^{\mu}$ and $\phi_{c}$ with $\alpha$ at large stresses suggests that the critical rigidity transition in shear thickening suspensions is simply a pre-cursor to the shear jamming transition.
@@ V Conclusions
A discrete element method (LF-DEM) based numerical simulations are carried out for systems of highly concentrated polydisperse shear thickening suspensions in 2D to understand the effects of particle size distribution on rigidity transition near shear jamming in the limit of large shear stress. The relative viscosity of different polydisperse suspensions is found to be identical to that of their statistically equivalent bidisperse systems. Furthermore, the viscosity divergence in 2D polydisperse and statistically equivalent bidisperse suspensions follow the Maron-Pierce model, similar to 3D systems. In the limit of large stress values, shear jamming phenomenon in polydisperse suspensions is found to be preceded by a critical rigidity transition similar to that observed by Santra et al. [51], in their study on bidisperse suspensions. Pebble game algorithm is implemented to evaluate an order-parameter from the fraction of rigid particles, $f_{\text{rig}}$, and its fluctuation, $\chi_{\text{rig}}$, which show identical variation with solid packing fraction for the polydisperse and statistically equivalent bidisperse suspensions. Such characteristics illustrate the similarity in the micro-structure of polydisperse and bidisperse suspensions. This further implies that the evolution of the flow micro-structure in shear thickening suspensions of polydisperse particles can be understood from the properties of relatively simple bidisperse systems.
@@
The critical transition observed in the polydisperse suspensions are characterized by determining the scaling behaviour of $f_{\text{rig}}$ and $\chi_{\text{rig}}$. The values of the scaling exponents $\beta=1/8$ and $\gamma=7/4$, for $f_{\text{rig}}$ and $\chi_{\text{rig}}$, respectively, are found to be consistent with the 2D-Ising model. Furthermore, the critical exponents $\gamma$ and $\nu$ are determined independently from finite size scaling analysis of the correlation length and susceptibility. In this context, we have used a protocol to compute the correlation length of the rigid clusters from the pair-correlation function of the rigid particles. Interestingly, from the data collapse the value of $\nu$ is found to be $1.64$, which is consistent with the value of exponent $\nu$ in 2D continuum percolation model. This essentially validates that the rigidity transition in shear thickening suspensions also corresponds to a percolation transition, which was qualitatively indicated in a previous study by van der Naald et al. [59]. Our investigation also indicates that the critical packing fraction $\phi_{c}$ is affected by the stiffness of the particles, moving the critical transition to higher packing fraction with decreasing spring constants. While only high stress values are considered in the present work to focus on the jamming transition well above the discontinuous shear thickening (DST) regime, flow behaviour and rigidity transition for polydisperse suspensions at lower stresses would be even more interesting because of the effect of CST-DST transition, which would be addressed in the future work. Nevertheless, the present study would be the basis to experimentally explore critical transition in shear thickening suspensions of more complex systems with heterogeneous particle size and shapes. Moreover, we show that both the bulk rheology and microscopic statistical features of polydisperse suspensions could be predicted from the behaviour of statistically equivalent bidisperse suspensions. We also demonstrate that rigidity transition in shear thickening suspensions with different particle size distributions have features which could be explained by equilibrium phase transition model, such as 2D Ising model, and percolation theory.

### v2

@@ IV.2 Finite size scaling
Figure 9: Finite size scaling of the correlation length $\xi$, susceptibility $\chi_{\text{rig}}$, and order-parameter $f_{\text{rig}}$ for polydisperse suspensions with number of particles $N=1000$, $2000$, $3000$ and $8000$, $\alpha=0.133$ at stress $\sigma=100$. Collapse of data for different system sizes are observed with scaling exponents $\beta\approx 0.14$, $\gamma\approx 2.4$ and $\nu\approx 1.33$. The insets display the corresponding unscaled data.
According to the present model, the frictional contacts between the particles depend on the spring constants for the normal and tangential components of the contact force. While in this study we have considered the particles to be non-deformable rigid spheres, the stiffness of the particles may still be correlated to the stiffness of the springs at the frictional contact. This section discusses the effects of polydispersity index and spring constants on the critical packing fraction and shear jamming. As presented in Fig. 10 (a), for a given value of stress ($\sigma$) and spring constant ($k_{n}$) the critical packing fraction $\phi_{c}$ varies non-monotonically with polydispersity index $\alpha$ and goes through a minima. The same trend is observed for the shear jamming fraction $\phi_{J}^{\mu}$, as shown in Fig. 10 (b). This is in contrast to the nature of variation of the jamming threshold ($\phi_{m}$) observed for frictionless polydisperse systems [15, 69], where $\phi_{m}$ is found to monotonically increase with polydispersity index ($\alpha$) for positive skewness ($s>0$) and show only week dependence on $\alpha$ for $s<0$. Notably, in the present system, while for the same value of $k_{n}$(=2000=2000) $\phi_{c}$ is smaller for the larger stress ($\sigma=100$) at all polydispersity indices (as shown in Fig. 10 (a)), $\phi_{J}^{\mu}$ shows a opposite behavior where it is higher for $\sigma=100$ as compared to $\sigma=25$. It is noteworthy that by proportionally lowering the value of $k_{n}$ to 500 for $\sigma=25$, both $\phi_{J}^{\mu}$ and $\phi_{c}$ are found to be higher for $\sigma=25$ as compared to $\sigma=100$ at all values of $\alpha$ (except for very small $\alpha$). Thus with decreasing spring stiffness, both $\phi_{J}^{\mu}$ and $\phi_{c}$ increases at any given stress suggesting the suspensions to become more flowable with softer particle contacts. This can be correlated to increasing flowability of softer deformable particles which could easily squeeze past each other under shear flow as compared to more rigid particles. Moreover, the similar variation of $\phi_{J}^{\mu}$ and $\phi_{c}$ with $\alpha$ at large stresses suggests that the critical rigidity transition in shear thickening suspensions is simply a pre-cursor to the shear jamming transition.
@@ V Conclusions
A discrete element method (LF-DEM) based numerical simulations are carried out for systems of highly concentrated polydisperse shear thickening suspensions in 2D to understand the effects of particle size distribution on rigidity transition near shear jamming in the limit of large shear stress. The relative viscosity of different polydisperse suspensions is found to be identical to that of their statistically equivalent bidisperse systems. Furthermore, the viscosity divergence in 2D polydisperse and statistically equivalent bidisperse suspensions follow the Maron-Pierce model, similar to 3D systems. In the limit of large stress values, shear jamming phenomenon in polydisperse suspensions is found to be preceded by a critical rigidity transition similar to that observed by Santra et al. [56], in their study on bidisperse suspensions. Pebble game algorithm is implemented to evaluate an order-parameter from the fraction of rigid particles, $f_{\text{rig}}$, and its fluctuation, $\chi_{\text{rig}}$, which show identical variation with solid packing fraction for the polydisperse and statistically equivalent bidisperse suspensions. Such characteristics illustrate the similarity in the micro-structure of polydisperse and bidisperse suspensions. This further implies that the evolution of the flow micro-structure in shear thickening suspensions of polydisperse particles can be understood from the properties of relatively simple bidisperse systems. One of the key findings from the present study suggests that the polydispersity only shifts the critical packing fraction $\phi_{c}$ without qualitatively changing the critical phenomenon, which could have been affected by the way particles pack with variation in polydispersity. However, within the limited scope of the present study with finite size systems we have not explicitly considered the effects of the nature of particle packing with polydispersity. To bring out any explicit dependence of the particle packing on the overall characteristics of the critical phenomenon, large scale simulations may be needed with larger system sizes. We may speculate that the nature of particle packing may appear as an external field variable in the scaling relations, similar to the role of a static friction coefficient investigated in another study [56], but it warrants detailed investigation.
@@
The critical transition observed in the polydisperse suspensions are characterized by determining the scaling behavior of $f_{\text{rig}}$ and $\chi_{\text{rig}}$. The values of the scaling exponents $\beta=0.164\pm 0.025$ and $\gamma\approx 2.4$, for $f_{\text{rig}}$ and $\chi_{\text{rig}}$, respectively, are found to be consistent with the 2D percolation transition. Furthermore, the values of the critical exponents $\beta$, $\gamma$ and $\nu$ are also approximated from the finite size scaling analysis of the correlation length, order-parameter and susceptibility. In this context, we have used a protocol to compute the correlation length of the rigid clusters from the pair-correlation function of the rigid particles. Notably, by choosing the values of $\beta\approx 0.14$, $\gamma\approx 2.4$ and $\nu\approx 1.33$, which are consistent with the values of critical exponents for standard 2D percolation transition, we have shown good data collapse within the framework of finite size scaling ansatz. This essentially validates that the rigidity transition in shear thickening suspensions also corresponds to a percolation transition, which was qualitatively indicated in a previous study by van der Naald et al. [66] for 2D suspensions, and quantitatively analyzed in an independent study by Goyal et al. [22] for 3D systems by computing percolation probability of 4-neighbor particles. Our investigation also indicates that the critical packing fraction $\phi_{c}$ is affected by the stiffness of the particles, moving the critical transition to higher packing fraction with decreasing spring constants. While only high stress values are considered in the present work to focus on the jamming transition well above the discontinuous shear thickening (DST) regime, flow behavior and rigidity transition for polydisperse suspensions at lower stresses would be even more interesting because of the effect of CST-DST transition, which would be addressed in the future work. Nevertheless, the present study would be the basis to experimentally explore critical transition in shear thickening suspensions of more complex systems with heterogeneous particle size and shapes. This is particularly important from the practical point of view since the ability to predict the critical volume fraction could prevent getting into undesirable jammed states in various industrial operations involving dense suspensions. Moreover, we show that both the bulk rheology and microscopic statistical features of polydisperse suspensions could be predicted from the behavior of statistically equivalent bidisperse suspensions. We also demonstrate that rigidity transition in shear thickening suspensions with different particle size distributions has features which could be explained by a continuous phase transition model, such as, 2D percolation theory.
Author Contributions
SKS and VT contributed equally by performing the computer simulations, collecting the data and developing analysis tools. SKS, VT and AS conceived and designed the analysis. All the authors performed the analysis and wrote the paper.
Conflicts of Interest
There are no conflicts to declare